# Diffusion Reward: Learning Rewards via Conditional Video Diffusion

### v2

@@ Abstract
Learning rewards from expert videos offers an affordable and effective solution to specify the intended behaviors for reinforcement learning tasks. In this work, we propose Diffusion Reward, a novel framework that learns rewards from expert videos via conditional video diffusion models for solving complex visual RL problems. Our key insight is that lower generative diversity is observed when conditioned on expert trajectories. Diffusion Reward is accordingly formalized by the negative of conditional entropy that encourages productive exploration of expert-like behaviors. We show the efficacy of our method over 10 robotic manipulation tasks from MetaWorld and Adroit with visual input and sparse reward. Moreover, Diffusion Reward could even solve unseen tasks successfully and effectively, largely surpassing baseline methods. Project page and code: diffusion-reward.github.io.
@@ 1 Introduction
Reward specification poses a fundamental challenge in reinforcement learning (RL), influencing the effectiveness and alignment of an agent’s learned behavior with the intended objectives. Manually designing dense reward functions is burdensome and sometimes impossible, particularly in real-world tasks such as robotic manipulation [31], where obtaining privileged state information is difficult. As a substitute, using sparse rewards is often favorable in these scenarios because of its low demand for manual effort [25]. Nonetheless, the sample efficiency of RL drops significantly due to insufficient supervision from sparse rewards.
@@
Figure 1: Overview. (top) We present a reward learning framework in RL using video diffusion models. We perform diffusion processes conditioned on historical frames to estimate conditional entropy as rewards to encourage RL exploration of expert-like behaviors. (bottom) The mean success rate of 10 visual robotic manipulation tasks demonstrates the effectiveness of our Diffusion Reward over 5 runs. Shaded areas are standard errors.
@@
Learning reward functions from expert videos offers a promising solution because of the low effort of video collection and dense task-execution information contained in the videos [5, 42]. Given unlabeled videos, generative models have been naturally investigated by researchers to extract informative rewards unsupervisedly for RL training [27, 35]. One classical approach builds on generative adversarial learning to learn a discriminative reward that discerns between agent and expert observations. While straightforward, these methods underutilize the temporal information, whose importance has been shown in solving RL [44], and performance is brittle to the adversarial training. In light of these issues, recent work leverages the VideoGPT [38] to encode temporal information, and directly use the predicted log-likelihood as rewards [9]. However, it struggles with modeling complex expert video distributions, particularly those with intricate dynamics. As shown in Figure 2, there exists a noticeable decline in the learned return is observed for out-of-distribution expert videos, despite sharing optimal behavioral patterns with in-distribution ones.
@@
Video diffusion models [17] — which have exhibited remarkable performance in the domain of computer vision — have shown great power in capturing the complex distribution of videos, such as text-to-video generation [16, 11] and video editing [4, 24]. Recent works have also examined their capability of modeling expert videos as generalizable planners in robotic manipulation tasks [8, 21]. Despite these advancements, extracting informative rewards from video diffusion models remains an understudied area, while shining great potential for guiding RL agents to acquire expertise from videos and generalizing to unseen tasks.
@@
In this work, we propose Diffusion Reward, a reward learning framework that leverages conditional video diffusion models to capture the expert video distribution and extract dense rewards for visual RL. Our key insight is that higher generative diversity is observed when conditioned on expert-unlike videos, while lower given expert videos. This rationale naturally instructs RL exploration on expert-like behaviors by seeking lower diversity. We therefore estimate entropy conditioned on historical frames, which formalizes our insight, and augment it with a novelty-seeking reward [3] and spare environment reward to form dense rewards for efficient RL. In addition, to accelerate the reward inference, we perform latent diffusion processes by utilizing vector-quantized codes  [12] for compressing high-dimensional observations.
@@
We empirically validate the efficacy of our framework through experiments on 10 visual robotic manipulation tasks, including 7 gripper manipulation tasks from MetaWorld [40] and 3 dexterous manipulation tasks from Adroit [28], exhibiting 38% and 35% performance improvements over the best-performing baselines given the same training steps, respectively. Furthermore, we surprisingly find that our learned reward can achieve fair zero-shot generalization performance on unseen tasks. Figure 1 overviews this work. The primary contributions of this work can be summarized as follows:
@@
  • We present Diffusion Reward, a novel reward learning framework that leverages the generative modeling capabilities of video diffusion models to provide dense reward signals for RL agents.
  • We show that our framework significantly outperforms baselines on 10 visual robotic manipulation tasks.
@@
  • We find that our pre-trained reward model could produce reasonable rewards and thus instruct RL on unseen tasks.
@@
Figure 2: Rewards from different video models. Results are averaged over 7 MetaWorld tasks with 10 random seeds for each task. Suboptimal represents videos with 25% of actions taken by random policy. VIPER (CE) and Diffusion Reward (LL) replace their original rewards with conditional entropy (CE) and Log-likelihood (LL), respectively. We observe that LL-based methods assign relatively low rewards to unseen expert videos while CE-based methods are able to assign near-optimal rewards to unseen expert videos. Moreover, such a boost is enhanced by the strong modeling ability of diffusion models.
@@ 2 Related Work
Learning rewards from videos. Extracting supervision signals from expert videos provides an affordable but effective solution for reward specification in RL [5, 29, 30]. A number of works have attempted to learn dense rewards indicating task progress by measuring the distance between current observation and goal image in the latent space [23, 42]. While promising, goal images are often hard to obtain out of simulation, limiting their applicability to open-world tasks [22]. On the contrary, generative models have been widely investigated to extract rewards unsupervisedly without future information. One representative approach builds on generative adversarial learning [35, 27] to discern expert-like and expert-like behaviors. Such an idea is further improved in [9] by predicting the log-likelihood as rewards, where the video prediction model is used to encode more temporal information. Nevertheless, it still struggles with modeling expert video distributions in complex tasks featuring intricate dynamics, thus being prone to produce unproductive rewards. In contrast, our methods leverage the powerful modeling abilities of diffusion models and estimate the negative of conditional as more discriminative rewards to expedite RL exploration. Moreover, our pre-trained reward could be generalized to unseen tasks better.
Figure 3: Video prediction results. Our video diffusion model could capture the distribution of expert videos from complex tasks. The outcomes are conditioned on two history frames and predictions. Ground truth has blue borders and prediction has orange borders.
Diffusion models for RL. Diffusion models have been widely investigated for RL to, for instance, improve the policy expresiveness [37, 14, 7], and augment experience [41, 6]. Apart from these, some works directly learn the diffusion models from offline data unconditional planners [19], or conditional planners specified by task returns [1]. The idea of conditional diffusion is further investigated in [8, 21] with text as task specification, where video diffusion models serve as planners associated with inverse dynamics. Unlike these methods, we intend to learn informative rewards via video diffusion models conditioned on historical frames to accelerate online RL. Such historical conditioning has been used in [18] to inform trajectory generation, which differs from our focus on reward learning as well. Our work is also close to Nuti et al. [26], which has also attempted to extract reward from two diffusion models that fit different behaviors in 2D tasks. Differently, we learn diffusion-based rewards simply from expert behaviors and achieve favorable performance on complex vision-based manipulation tasks.
@@ 3 Preliminaries
Problem formulation.
We consider an RL agent that interacts with the environment modeled as a finite-horizon Markov Decision Process (MDP), which is defined by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}$ is the environment transition function, $\mathcal{R}$ is the reward function, and $\gamma$ is the discount factor. The goal of the RL agent is to learn an optimal policy $\pi$ that maximizes the expected return $\mathbb{E}[\sum_{k=0}^{K-1}\gamma^{k}r_{k}]$.
In this work, we focus more specifically on high-dimensional state space with a binary sparse (i.e., 0/1 reward as a success indicator of the task). Particularly, we consider RGB images $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ as the state observed by the agent. This setting is motivated by the real-world application of RL such as robotics, where vision-based sensory data is more available and specifying sophisticated reward often requires tedious hand-engineering, sometimes even intractable. However, this poses great difficulty to an RL agent, as a large amount of interaction with environments is often required, known as sample efficiency.
Expert videos. To improve the sample efficiency of RL, we assume a set of unlabeled videos generated by the expert policies are accessible by the agent. Notably, these videos are action-free and gathered from multiple tasks without task identification. We denote it as $\mathcal{D}=\{\mathcal{D}^{1},\mathcal{D}^{2},...,\mathcal{D}^{N}\}$, where $\mathcal{D}^{i}$ is the demonstrated videos from task $i$. Each set of demonstrated videos $\mathcal{D}^{i}$ contains multiple expert trajectories $\tau=\{\bm{x}_{0},\bm{x}_{1},...,\bm{x}_{K-1}\}\in\mathcal{S}^{K}$. Our goal now turns to learning effective reward functions of such videos to accelerate the online exploration of RL agents.
@@
Table 1: Quantitative comparison of video models. Results are evaluated on real robot videos and simulation videos from 7 MetaWorld tasks, demonstrating that video diffusion model generates videos of higher quality than VideoGPT.
4 Method
We introduce Diffusion Reward, a novel framework that learns rewards from expert videos via conditional video diffusion models for solving downstream visual RL problems, as illustrated in Figure 1. At a high level, our method leverages entropy information from video diffusion models pre-trained on expert videos to encourage RL agents to explore expert-like trajectories more. In Section 4.1, we first model the expert videos with the conditional video diffusion model in the latent space. In Section 4.2, we formalize our key insight of Diffusion Reward by estimating the history-conditioned entropy and using its negative as reward signals for RL training. We summarize our framework in Algorithm 1 and present the overview and key implementation details of our framework below.
4.1 Expert Video Modeling via Diffusion Model
Diffusion models [32] are probabilistic models that aim to model the data distribution by gradually denoising a normal distribution through a reverse diffusion process [15]. These models showcase their power in capturing highly complex distributions and generating samples that exhibit intricate dynamics, motion, and behaviors in RL literature [19, 1]. Unlike prior works that model expert videos as planners, we aim to learn reward functions from the diffusion model trained on expert videos for RL. This motivates our video models to achieve fast inference speed and encode temporal information.
Figure 4: Conditional entropy as rewards. (left) Aggregated learned rewards over 7 tasks from MetaWorld, which assigns expert policy high rewards while random policy low rewards. (right) Normalized conditional entropy on both seen and unseen tasks from MetaWorld, demonstrating that our rewards can distinguish trajectories of high quality from those of decreasing qualities and indicating its potential generalization ability. (top) One pair of the evaluated expert and random trajectories is exemplified above.
Latent diffusion process. Specifically, we first train an encoder unsupervisedly from expert videos to compress the high-dimensional observations. Here we use the VQ-GAN method [10] to represent the image $\bm{x}$ with a vector-quantized latent code $\bm{z}=Q(E(\bm{x}))$, where $E$ is the encoder and $Q$ is the element-wise quantizer. The whole video is then represented by a sequence of latent variables $\tau=\{\bm{z}_{0},\bm{z}_{1},...,\bm{z}_{K-1}\}$, where we overwrite the definition of $\tau$ without ambiguity. Subsequently, the forward process applies noise $\epsilon$ in the latent space at each time step $t\in 0,...,T$ to the data distribution $\bm{z}_{k}$, resulting in a noisy sample $\bm{z}_{k}^{t}$, where $\bm{z}_{k}^{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{k}^{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, and $\bar{\alpha}$ is the accumulation of the noise schedule over past timesteps. To fit data distribution, we learn a parameterized variant of noise predictor $\bm{\epsilon}_{\theta}(\bm{z}_{k}^{t})$ that aims to predict the noise $\bm{\epsilon}$ during the forward process. Then the parameterized reverse process $p_{\theta}(\bm{z}_{k}^{t-1}|\bm{z}_{k}^{t})$ can be approximated and performed by iterative denoising a initial distribution.
Historical frames as condition. To utilize the temporal information from expert videos with the power of video diffusion, we further condition the reverse process with history frames, i.e., $p_{\theta}(\bm{z}_{k}^{t-1}|\bm{z}_{k}^{t},\bm{z}_{c})$, where $\bm{z}_{c}$ is the concatenation of all historical frames $[\bm{z}_{0},...,\bm{z}_{k-1}]$. This can also be viewed as matching the distribution of expert and agent trajectories [9]. In practice, we use a subsequence of historical frames as a condition to ensure higher computation efficiency while maintaining temporal information. Subsequently, one can perform the reverse process from a randomly sampled noise to generate the latent code of future frames and decode the code for video prediction, as shown in Figure 3. In this work, we use VQ-Diffusion [12] as our choice of video diffusion model due to its good performance and compatibility with vector-quantized latent code, but our framework can in principle adopt any off-the-shelf video diffusion models. We first tokenize each latent code $\bm{z}$ indexed by its indices that specify the respective entry in the learned codebook of VQ-GAN, and take as a condition embedding the concatenated tokens. The embedding is then fed to the decoder that contains 16 transformer blocks and softmax layers with cross attention.
4.2 Conditional Entropy as Rewards
While previous studies have explored the use of log-likelihood as rewards with video prediction models, exemplified by works such as VIPER [9], this approach encounters two primary challenges. Firstly, it struggles with accurately modeling the distribution of complex expert videos featuring intricate dynamics, as shown in Table 1. Secondly, the moderate video modeling ability leads to undesired rewards. This issue is evident in Figure 2, where a significant drop in learned rewards between in-distribution expert videos and out-of-distribution ones, though both sets of videos demonstrate optimal behaviors.
Key insight behind Diffusion Reward. We address these challenges by harnessing the great generative capability of video diffusion models. Our observations indicate increased generation diversity with unseen historical observations (rand.) and reduced diversity with seen ones (expert), as shown in Table 2. This gives rise to the key insight of our proposed Diffusion Reward: diffusion conditioned on expert-like trajectories exhibits lower diversity where the agent ought to be rewarded more, and the opposite holds on unexpert-unlike ones. Such diversity discrimination not only incentivizes RL agents to chase expert-like behaviors but also enhances exploration through the stochasticity induced by the diffusion process.
@@ 4.2 Conditional Entropy as Rewards
Table 2: Quantitative results of generative diversities. Referring to [45], generative diversity is proportional to distances among a batch of generated videos. Expert trajectories, which are seen during training, show the lowest diversity while the unseen random ones exhibit the highest diversity. Results are over 7 tasks from MetaWorld and trajectories are generated by $\epsilon$-greedy policies.
Algorithm 1 Diffusion Reward

### v3

# diffusion-reward.github.io

@@ Abstract
Learning rewards from expert videos offers an affordable and effective solution to specify the intended behaviors for reinforcement learning (RL) tasks. In this work, we propose Diffusion Reward, a novel framework that learns rewards from expert videos via conditional video diffusion models for solving complex visual RL problems. Our key insight is that lower generative diversity is exhibited when conditioning diffusion on expert trajectories. Diffusion Reward is accordingly formalized by the negative of conditional entropy that encourages productive exploration of expert behaviors. We show the efficacy of our method over robotic manipulation tasks in both simulation platforms and the real world with visual input. Moreover, Diffusion Reward can even solve unseen tasks successfully and effectively, largely surpassing baseline methods. Project page and code: diffusion-reward.github.io.
Keywords:
Reward Learning from Videos Robotic Manipulation Visual Reinforcement Learning
@@ 1 Introduction
Reward specification is a fundamental challenge in reinforcement learning (RL), determining the effectiveness and alignment of an agent’s learned behavior with the intended objectives. Manually designing dense reward functions is both labor-intensive and sometimes infeasible, particularly in real-world tasks such as robotic manipulation [41], where obtaining privileged state information is difficult. As an alternative, using sparse rewards is often favorable in these scenarios because of its low requirement for manual effort [30]. Nonetheless, the insufficient supervision provided by sparse rewards hinders the sample efficiency of RL, potentially imposing an unaffordable burden on data collection.
@@
Figure 1: Overview of Diffusion Reward. (left) We present a reward learning framework in RL using video diffusion models pretrained with expert videos. We perform diffusion processes conditioned on historical frames to estimate conditional entropy as rewards to encourage RL exploration of expert-like behaviors. (right) The mean success rate of 10 visual robotic manipulation tasks demonstrates the effectiveness of our proposed Diffusion Reward over 5 runs. Shaded areas are standard errors.
@@
Learning reward functions from expert videos presents a promising solution because of the low effort of video collection and dense task-execution information contained in the videos [6, 55]. Among these, generative models have been naturally investigated by researchers to extract informative rewards unsupervisedly for RL training [33, 46]. One classical approach builds on generative adversarial learning to learn a discriminative reward that distinguishes between agent and expert behaviors. While straightforward, these methods underutilize the temporal information, whose importance has been shown in solving RL [57], and performance is brittle to the adversarial training. In light of these issues, recent work pre-trains VideoGPT [50] to encode temporal information, and directly use the predicted log-likelihood as rewards [10]. However, they encounter difficulties in modeling complex expert video distributions and accordingly the proposed rewards, especially in tasks with intricate dynamics. As shown in Figure 3, a discernible decline in the learned return is evident for out-of-distribution expert videos, despite their alignment with optimal behaviors in in-distribution videos.
@@
Video diffusion models [17] — which have exhibited remarkable performance in the domain of computer vision — have shown great power in capturing the complex distribution of videos, such as text-to-video generation [15, 11] and video editing [5, 27]. Recent studies have also investigated their potential in modeling expert videos to serve as generalizable planners in robotic manipulation tasks [19, 9, 21]. Despite these advancements, extracting informative rewards from video diffusion models remains an understudied area, while shining great prospects in guiding RL to acquire expertise from videos and generalize to unseen tasks.
@@
In this work, we propose Diffusion Reward, a reward learning framework that leverages video diffusion models to capture expert video distribution and pre-train dense reward function for visual RL. Our key insight is that diffusion conditioned on expert-like trajectories exhibits lower generative diversity than that conditioned on expert-unlike ones. This rationale naturally instructs RL exploration on expert-like behaviors by seeking low generative diversity. We accordingly estimate entropy conditioned on historical frames, which formalizes our insight, and augment it with a novelty-seeking reward [4] and spare environment reward to form dense rewards for efficient RL. In addition, to accelerate the reward inference, we perform latent diffusion processes by utilizing vector-quantized codes  [13] for compressing high-dimensional observations.
@@
We empirically validate the effectiveness of our framework through experiments on 10 visual robotic manipulation tasks, including 7 gripper manipulation tasks from MetaWorld [52] and 3 dexterous manipulation tasks from Adroit [36], exhibiting 38% and 35% performance improvements over the best-performing baselines, respectively. Additionally, we surprisingly observe favorable zero-shot generalization performance of our learned reward on unseen tasks. Furthermore, reasonable rewards produced on real robot tasks and associated favorable offline RL performance demonstrate the real-world applicability of Diffusion Reward. Figure 1 overviews this work. The primary contributions of this work can be summarized as follows:
@@
  • We introduce Diffusion Reward, a novel reward learning framework that leverages the generative modeling capabilities of video diffusion models to provide informative dense reward signals for efficient RL.
  • We demonstrate significant outperformance of our framework compared to baselines across robotic manipulation tasks in simulation and the real world.
@@
  • We observe that our pretrained reward model can instruct RL on unseen tasks and produce reasonable rewards on real robot tasks.
@@
Figure 2: Analysis of video models and rewards. VIPER (CE) and Diffusion Reward (LL) replace original rewards with conditional entropy (CE) and log-likelihood (LL), respectively. CE-based rewards assign near-optimal rewards to unseen expert videos. Such a boost is enhanced by the strong modeling ability of diffusion models. Results are averaged over 7 MetaWorld tasks. Suboptimal represents videos with 25% randomly-taken actions.
Figure 3: Reward analysis. (left) Our learned rewards are aggregately higher for expert behaviors over 7 tasks from MetaWorld. (right) Varied-quality trajectories can be distinguished from seen and unseen tasks. (top) Examples of trajectories.
@@ 2 Related Work
Learning rewards from videos.
Extracting rewards from expert videos provides an affordable but effective solution for reward specification in RL [6, 39, 38]. Several studies have endeavored to learn dense rewards that indicate task progress by measuring the distance between current observation and goal image in the latent space [26, 55, 3]. While promising, obtaining goal images out of simulation can be challenging, thus limiting their applicability to open-world tasks [25]. In contrast, generative models have been widely explored to extract rewards without access to future information. One representative approach, belonging to state-only imitation learning[45, 35, 59], builds on generative adversarial learning [46, 33] to distinguish expert-like and expert-unlike behaviors with an online paradigm. This idea is further refined in [10], wherein a reward function is pretrained from videos, and the log-likelihood is predicted as rewards. Nonetheless, it encounters challenges in accurately modeling expert video distributions in complex tasks featuring intricate dynamics, thereby rendering them susceptible to generating unproductive rewards. In contrast, our methods leverage the powerful modeling abilities of diffusion models and estimate the negative of conditional as more discriminative rewards to expedite RL exploration.
Diffusion models for RL.
Diffusion models have been widely investigated for RL to, for instance, improve the policy expresiveness [48, 14, 8], and augment experience [53, 7]. Apart from these, some works directly learn the diffusion models from offline data unconditional planners [19], or conditional planners specified by task returns [2]. The idea of conditional diffusion is further investigated in [9, 21] with text as task specification, where video diffusion models serve as planners associated with inverse dynamics. Unlike these methods, we intend to learn informative rewards via video diffusion models conditioned on historical frames to accelerate online RL. Such historical conditioning has been used in [18] to inform trajectory generation, which differs from our focus on reward learning as well. Our work is also close to [31], which extracts rewards from two diffusion models that fit different behaviors in state-based tasks. Differently, we pretrain reward models without action labels and additional task-agnostic data.
Pretraining for RL.
Numerous methods have been explored for directly pretraining effective policies in RL from data. These methods include those employing offline RL techniques before online finetuning [28, 23, 22] as well as approaches learning behavioral (skill) priors [34, 40, 37]. However, they require the additional collection of rewards, action labels, or task-agnostic data, which poses certain sacrifices. In contrast, our approach focuses on pretraining rewards solely from action-free expert-only videos, alleviating the need for such additional data sources. Some studies have also explored pretraining visual representations using datasets similar to ours [49, 29, 32, 54]. Nevertheless, these approaches do not extract dense task-execution signals to guide RL.
@@ 3 Preliminaries
Table 1: Quantitative comparison of video models. Results are evaluated on real dexterous robotic manipulation videos and simulation videos from 7 MetaWorld tasks, demonstrating that the video diffusion model generates videos of higher quality than VideoGPT.
@@
Table 2: Generative diversities. Expert trajectories over 7 MetaWorld tasks, seen during training, show the lowest generative diversity, which is proportional to distances among generated videos [58].
Problem formulation.
We consider an RL agent that interacts with the environment modeled as a finite-horizon Markov Decision Process (MDP), which is defined by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}$ is the environment transition function, $\mathcal{R}$ is the reward function, and $\gamma$ is the discount factor. The goal of the RL agent is to learn an optimal policy $\pi$ that maximizes the expected return $\mathbb{E}[\sum_{k=0}^{K-1}\gamma^{k}r_{k}]$ over $K$ timesteps.
In this work, we focus more specifically on high-dimensional state space with a binary sparse (i.e., 0/1 reward as a success indicator of the task). Particularly, we consider RGB images $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ as the state observed by the agent. This setting is motivated by the real-world application of RL such as robotics, where vision-based sensory data is more available and specifying sophisticated reward often requires tedious hand-engineering, sometimes even infeasible. However, this poses great difficulty to an RL agent, as a large amount of interaction with environments is often required, known as sample efficiency.
Expert videos.
To improve the sample efficiency of RL, we assume a set of unlabeled videos generated by the expert policies are accessible by the agent. Notably, these videos are action-free and gathered from multiple tasks without task identification. We denote it as $\mathcal{D}=\{\mathcal{D}^{1},\mathcal{D}^{2},...,\mathcal{D}^{N}\}$, where $\mathcal{D}^{i}$ is the demonstrated video set of task $i$. Each video set $\mathcal{D}^{i}$ consists of multiple expert videos $\tau=\{\bm{x}_{0},\bm{x}_{1},...,\bm{x}_{K-1}\}\in\mathcal{S}^{K}$. Our goal is to pretrain effective reward functions from expert videos to accelerate the online exploration of RL. We present our reward learning framework in the following section.
4 Method
We introduce Diffusion Reward, a novel framework that learns rewards from expert videos via conditional video diffusion for solving downstream visual RL problems, as illustrated in Figure 1 and Algorithm 1. In Section 4.1, we first model the expert videos with the video diffusion model in the latent space. In Section 4.2, we formalize our key insight of Diffusion Reward by estimating the history-conditioned entropy and using its negative as rewards for RL.
4.1 Expert Video Modeling via Diffusion Model
Diffusion models [42] are probabilistic models that aim to model the data distribution by gradually denoising an initial distribution through a reverse diffusion process [16]. These models showcase their power in capturing highly complex distributions and generating samples that exhibit intricate dynamics, motion, and behaviors in RL literature [19, 2]. Unlike prior works that model expert videos as planners, we aim to learn reward functions from the diffusion model trained on expert videos for RL. This motivates our video models to achieve fast inference speed and encode sufficient temporal information.
Algorithm 1 Diffusion Reward for online RL